Other beamlines: 8.3.1 12.3.1

8.3.1 last updated: Wed Nov 26 22:41:59 PST 2025
sample: no beam mark

camera:

hutch:

no movement for 04:34:30 since Nov 26 18:07:33 2025
beam:


hutch is dark
feedback system is on

last tuneup reached 287% of max at Nov 26 10:30:28 2025
final diode current was 20204.3 nA, equivalent to 6.126e+10 photons/s at div=0.6x0.1 mrad, 11145.0 eV and 500 mA (4.83 kGy/s)

cryostream is off at 294.67K, nozzle out, lN2 level is at 27% and set to autofill
expected to reach 20% at Thu Nov 27 04:46:33 PST 2025

CHL robot dewar is dry, level: unknown% and idle, supply idle at 0.98 psi, as of 49 s ago.

current user is AVAILABLE (until 8pm) (on-site with 01:17:57 left
next user is Maintenance at Nov 27 00:00:00
door closed for 04:34:31 since Nov 26 18:07:17 2025

00:00:08 ago last control activity at Nov 26 22:41:48 2025
05:24:03 ago at Wed Nov 26 17:17:53 PST 2025
collected: /data/mcfuser/ucsf/fraser/25Nov2025/mac1/mac1_19/mac1_19_avi3731_bg_2_00719.cbf

Exposure_time 0.1990500
Wavelength 1.11243
Detector_distance 175.4
Start_angle 718.0000
beam stop:

last 10 beam stops:



















/data is 86% full, still room for 1384242 images.

Other beamlines: 8.3.1 12.3.1